OpenStudy (anonymous):

What is the net force and acceleration of a ball of mass 0.750 kilograms that is subjected to forces of 11.00 newtons and 7.00 newtons from opposite directions?

OpenStudy (anonymous):

Net force on ball Fnet = F1 + F2 = 11 + (-7) = 4N Fnet=Ma a=Fnet/M = 4/0.75 = 16/3 N = 5.33 N